OpenStudy (jocelyn):

when 36 is subtracted from the square of a number, the result is five times the number. What is the positive solution? (1) 9 (2) 6 (3) 3 (4) 4

OpenStudy (anonymous):

Set up an equation: x^2 - 36 = 5x Now plug in the answers they give you and see which one comes out correctly. 9^2 - 36 = 5 x 9 81-36 = 5 x 9 45 = 5 x 9 45 = 45 9 works, so it is the answer.